# A geometric phase approach to quark confinement from stochastic gauge-geometry flows

### v1

@@ V Wilsonian approach to QCD and topology changes
As a second point, we have to consider the (potential) convergence behaviour of the RF. The behaviour depends strongly on the topology of the 3-manifold and its underlying geometric structure. In Appendix A we describe the eight possible geometries on 3-manifold and the decomposition of the 3-manifold into geometric pieces. For now, we will state the general behaviour: every spherical geometry has a finite extinction form of the RF; furthermore, the other geometries except of the hyperbolic geometry leading to a pinching of the underlying 3-manifold. Therefore, only the RF of the hyperbolic geometry can be written as a stochastic equation away from the fixed point.
VI Proposal for QCD confinement
@@
Geometric fluctuations due to the stochastic RF induce changes in the field lines following the pattern of knot concordance described in Section V. In this setting, the expectation of the knotted and linked Wilson loops is computed by considering the vev of the product of two Wilson loops as
$\langle\mathcal{W}_{\gamma_{1}}(A)\mathcal{W}_{\gamma_{2}}(A)\rangle=$ $\frac{1}{Z}\int\mathcal{D}A\,\mathcal{W}_{\gamma_{1}}(A)\,\mathcal{W}_{\gamma_{2}}(A)$ (46)
@@
where $S_{\rm YM}$ indicates the Yang-Mills action, and the normalization constant $Z$ is the partition function. Following the analogy of the Chern-Simons case in Cotta-Ramusino et al. (1990), we expect that $\langle\mathcal{W}_{\gamma_{1}}(A)\mathcal{W}_{\gamma_{2}}(A)\rangle$ depends on the linking of $\gamma_{1}$ and $\gamma_{2}$. For Chern-Simons, it is known that switching crossings produces certain skein relations, as shown in Ref. Cotta-Ramusino et al. (1990), which is the reason why the vev of Wilson loops gives link invariants such as the Jones polynomial. In the case of Yang-Mills, we do not expect the result to produce link invariants, but we conjecture that the linking results in interactions that influence confinement. Since linking is a property induced by stochastic fluctuations, as suggested in this article following the stochastic RF, it follows that confinement properties are induced by geometric properties. While we defer a detailed study of Eq. (46) and its relation to confinement to subsequent studies, we emphasize that one can compute $\langle W_{\gamma_{1}}(A)W_{\gamma_{2}}(A)\rangle$ also in the case where $\gamma_{2}$ is obtained from $\gamma_{1}$ by a small shift as in the case of the self-linking number. This corresponds to the presence of self-interaction terms.
VII Conclusions and outlook
We outlined the proposal for a novel geometric phase approach to confinement in Yang-Mills theories. The crucial aspects of the framework we have developed can be summarised as follows:
@@ VII Conclusions and outlook
  • ii) the gravitational back-reaction to the Yang-Mills fields induces stochastic fluctuations of topology of the sub-manifolds that represent matter;
  • iii) in dual picture, changes of topology corresponds to the generation of gauge vortices, which finally source the geometric phase of hadrons’ ground states.
@@
First of all, the formation at the fermi scale of a non-vanishing condensate is eminently related to the topological properties of the manifold. But the energy scales that defines the size of the ground states is still related to dimensional transmutation. Nonetheless, it might be tempting in this framework to relate the energy scale of confinement to amplitude of the stochastic noises that enter the Einstein–Yang-Mills stochastic flow. In Lulli et al. (2022) the amplitude of the gravitational stochastic noise was traced back to the variance of its fluctuations, and shown to be proportional to cosmological constant. On the other hand, for the gauge fields it might be natural to relate the origin of the amplitude of the stochastic noise to chaos effects in the dynamics of the chromo-electric and chromo-magnetic fluxes.
@@
Dimensional transmutation can be also reconsidered in the framework we are proposing. We may indeed account for the breakdown of the conformal symmetry at the quantum level, as being induced by the stochastic geometric flow, while the system is out of equilibrium.
The stochastic time regulates the out-of-equilibrium dynamics, making evident that the breakdown of the conformal symmetry (at the quantum level) emerges at non-zero stochastic time distances, and out of equilibrium. Scales can be then individuated in the theory through the coupling between gravity and matter. It is indeed the Newton constant $G$ that provides a dimension-full constant to introduce new physical scale. On the other hand, it is the interplay between the coupling constant of the theory of gravity and the specific coupling constant of matter that generate the physical relevant scale for the QFT under investigation.
Two specific examples of emergence of physical scales through dimensional transmutation – induced by the coupling with matter, and reminiscent of the conformal anomaly as sourced by the trace of the energy-momentum tensor – are provided by the theory of QED (the Bohr radius of the electron, that regulates the “size” of the Hydrogen atom) and the theory of QCD (the $\Lambda_{\rm QCD}$ scale). In these two cases, the physical scales should be provided by a running, induced by the stochastic geometric flow, of the values of $G$, together with a rescaling with the respect to $g_{\rm em}$ and $g_{\rm s}$ of the gravitational coupling constant $G$. Hence, it would be tempting to consider
@@
for the contribution from the QCD sector $\delta^{\rm QCD}\Lambda$ to the cosmological constant $\Lambda$.
The so-called “Vacuum Catastrophe” reflects the fundamental fact that the energy density of the quark-gluon condensate responsible for the color confinement in QCD is off by over forty orders of magnitude and has a wrong sign compared the observable cosmological $\Lambda$-term value. In the Yang-Mills effective action approach Savvidy (1977); Matinyan and Savvidy (1978) it is possible to show that the exact compensation of the electric (metastable) and magnetic (stable) gluon condensate components localised into vacuum domains spatially separated by domain walls in the QCD vacuum (averaged over space-time volumes above the fermi scale) is the necessary and sufficient condition for confinement in QCD Addazi et al. (2019a, b). Provided that the energy scales of “electric gluon” and “magnetic gluon” condensation are not the same, the latter are formed at different space-time separations. However, they evolve towards the same values of the energy density, but with opposite signs, due to their cosmological attractor nature. This effectively causes the cancellation of “electric gluon” and “magnetic gluon” contributions to the QCD ground state in macroscopic volumes, i.e. in the deep infrared limit of the theory Pasechnik et al. (2013a); Addazi et al. (2022). The gravitational interactions then break the discrete mirror symmetry of the QCD vacuum already in the semi-classical approximation. Such a breaking causes a non-zeroth (positive) gravitational correction to the resulting effective QCD-induced cosmological constant $\delta^{\rm QCD}\Lambda$ found to be close (within an order of magnitude) to the observed value Pasechnik et al. (2013b).
Such a cancellation emerges in experiment as a complete disappearance of the colored degrees of freedom from the theory in the IR limit, i.e. with characteristic wave lengths beyond the fermi scale, as they are “trapped” within the domains at the QCD length scale. This is fully compatible with the classical limit of the Yang-Mills theory where the conformal anomaly dynamically vanishes and where only a matter-like (non-relativistic) QCD medium (hadron gas) remains at large separations. For the exact cancellation and, hence, for color confinement the restoration of a discrete (mirror) symmetry between the “electric gluon” and “magnetic gluon” contributions at the level of the effective Lagrangian is an intrinsic property of pure Yang-Mills theory and the RG flow equations.
At the next stage, it would be mandatory to interconnect between the effective Yang-Mills action approach, featuring the dimensional transmutation and an emergence of the mirror symmetry of the Yang-Mills vacua, the respective RG flow and the RF formalism.

### v2

@@ VI Wilsonian approach to QCD and topology changes by the stochastic geometry flow
As a second point, we have to consider the (potential) convergence behavior of the geometry flow. The behavior depends strongly on the topology of the 3-manifold and its underlying geometric structure. In Appendix A we describe the eight possible geometries on 3-manifold and the decomposition of the 3-manifold into geometric pieces. We only state here the general behavior: every spherical geometry has a finite convergent time in the geometry flow — except for the hyperbolic geometry, the other geometries lead to a pinching of the underlying 3-manifold. Therefore, only the geometry flow for the hyperbolic geometry can be written as a stochastic equation away from the fixed point.
We may now go a step further. We discussed in section IV the pairwise appearance of a Wilson knot and a t’Hooft loop. This configuration has a simple interpretation. Let us focus on the trivial case, in which one Wilson loop and one t’Hooft loop produce the 3-sphere. This decomposition is known as Heegard decomposition of the 3-sphere $S^{3}=(D^{2}\times S^{1})\cup(S^{1}\times D^{2})$. One can understand this decomposition by a simple argument. The 4-disk $D^{4}$ has the 3-sphere as boundary $\partial D^{4}=S^{3}$. We then use the relation $D^{4}=D^{2}\times D^{2}$ and build the boundary $S^{3}=\partial D^{4}=\partial(D^{2}\times D^{2})$, to derive $\partial(D^{2}\times D^{2})=(\partial D^{2}\times D^{2})\cup(D^{2}\times\partial D^{2})$, leading to the final result.
For the case in which a loop (or unknot) is changed to a knot, then the decomposition can be also applied. Nonetheless, one obtains a homology 3-sphere, i.e. a 3-manifold with the same homology than a 3-sphere. This procedure is known as ±$\pm 1$ Dehn surgery along a knot. Then the change of the loop to a knot can be also expressed as a cobordism from the 3-sphere to a homology 3-sphere.
We are now able to determine the scaling behavior of the Wilson functional. At this purpose, we use the geometry of the knot complement. As previously discussed, most knot complements admit a hyperbolic geometry. Then the surface that is bounded by this knot (the Seifert surface) has a fixed area (determined by Mostow rigidity).
Furthermore, the knot complement has a geometry of constant curvature, which affects also the curvature of the Wilson knot. We then derive the desired behavior
$\intop_{S}F=\intop_{S}d(g_{a}^{-1}dg_{a})\sim{\rm Area}(S)$
for large fluctuations or large curvatures. For small fluctuations one cannot find such a surface, and the integral scales only w.r.t. length of the loop.
VII Scaling formula for the topology change
In this section, we discuss the fluctuations of the geometry flow, as expressed in terms of the fluctuations of the Wilson loop that lead to the topology change. As we discussed earlier, the stochastic geometry flow is connected with Einsteins equation. Thus the energy scale to be involved must be the Planck scale. One should then require a very high value of the energy, naturally the Planck energy, to produce a real change in the system. We shall focus on the direct consequences of the topology change as induced by the fluctuations of the Wilson loop.
The stochastic geometry flow leads to fluctuations of the Wilson loop. We allow for the topology change by using a cobordism, i.e. a 2-manifold connecting the unknotted Wilson loop into a knotted Wilson loop. This change is denoted as (ribbon) concordance. We may now ask ourselves what is the simplest topology change. Here, we need a 4-dimensional perspective to obtain an answer. We previously described the configuration in which a Wilson loop is linked to a t’Hooft loop as a 3-sphere. A change from a unknotted Wilson loop to a knotted Wilson loop is equivalent to a change from the 3-sphere to a homology 3-sphere. This change can be also described by a cobordism, known as homology cobordism. In the simplest case, one needs a pair of one 1-handle and one 2-handle to induce the simplest topology change. We now have to face the problem how to derive a scale for the homology 3-sphere from the scale of the 3-sphere (with Planck scale). We first focus on gaining a conceptual understanding of this question, before proceeding to the calculation of the concrete value. The approach was partly developed in Asselmeyer-Maluga and Krol (2018) where it was used to calculate the cosmological constant.
The change of the topology from the 3-sphere $S^{3}$ to some homology 3-sphere $\Sigma$ is described by a smooth 4-manifold $W$ with boundary $\partial W=S^{3}\sqcup\Sigma$, the so-called (homology) cobordism. Customarily, this change is described in the reversed way, i.e. the 1-handle and the 2-handle have to cancel each other. This process needs an embedded disk, which is complicated to be constructed in 4D. Instead, we need an infinite tower $CH$ of disks arranged along an infinite tree $T_{CH}$, usually called a Casson handle $CH$. The infinite tree representing this Casson handle must be embedded in a compact subset, which is only possible by using a subset with hyperbolic metric. For the embedding of the tree, it is enough to use a 2D model, i.e. the hyperbolic space $\mathbb{H}^{2}$. There are many isometric models of $\mathbb{H}^{2}$. We will use later the Poincaré disk model, but for our current purpose it suffices to deploy the half-plane model, selecting the hyperbolic metric
$ds^{2}=\frac{dx^{2}+dy^{2}}{y^{2}}$ (54)
in order to simplify the calculations. It is anyway possible to switch between the two isometric models, namely the Poincaré disk model and the half-space model.
The infinite tree $T_{CH}$ is embedded along the $y-$axis, so we may set $dx=0$. The tree $T_{CH}$, as the representative for the Casson handle $CH$, can be seen as metric space instead of a simplicial tree. While focusing on a simplicial tree, one is only interested in the structure given by the number of levels and branches. The tree $T_{CH}$, as a metric space (the so-called $\mathbb{R}-$tree), has the property that any two points are joined by a unique arc isometric to an interval in $\mathbb{R}$. Then the embedding of $T_{CH}$ is provided by the identification of the coordinate $y$ with the coordinate of the tree $a_{T}$ representing the distance from the root. This coordinate is a real number, and we can build the new distance function after the embedding as
$ds_{T}^{2}=\frac{da_{T}^{2}}{a_{T}^{2}}\,.$
As we have previously discussed, the tree $T_{CH}$ grows with respect to a time parameter, so that we need to introduce an independent “time scale” $t$. From the physics point of view, this time scale describes the partition of the tree into slices. Then a natural choice seems to be to set
$ds_{T}^{2}\sim dt^{2}\,.$
Here the time scale is related to the hyperbolic distance
$\frac{da_{T}^{2}}{a_{T}^{2}}=\frac{1}{L^{2}}dt^{2}$ (55)
via the scale $L$ of the hyperbolic structure for the 3-manifold $Y_{n}$ (see below) with $n=[a_{T}]$. However this equality is only a heuristic argument. The 3-manifolds $Y_{n}$ are intermediate steps in the topology changing processes induced by the corresponding level in the Casson handle, constrained by the values $Y_{0}=S^{3}$ and $Y_{\infty}=\Sigma$. A rigorous mathematical argumentation for the embedding of the Casson handle can be found in the Appendix B.
In the Casson handle $CH$, we have an infinite tower $CH=\bigcup_{n}T_{n}$ of disks (with self-intersections) arranged in levels $T_{n}$. For every level $T_{n}$, one has a scaling $L$. We then choose the time $t$ to be equal to the level $n=[t]$ for the integer part — remember the tree is a metric tree. Therefore, we are forced to make the identification
$Re(\xi)=\frac{t}{L}$
within the quadratic differential, defining the tree $T_{CH}$ — see the Appendix B. Then the length in the embedded tree is given by $d(Re(\xi$)), the measure of the vertical foliation. Now, the growth $ds_{T}^{2}$ of the tree with respect to the hyperbolic structure is provided by the measure $d(Re(\xi))^{2}$ of the vertical foliation, or
$ds_{T}^{2}=\frac{da_{T}^{2}}{a_{T^{2}}}=d\left(\frac{t}{L}\right)^{2}\,,$
which is in agreement with previous heuristic, dimensional argument. This equation can be formally integrated yielding the expression
$a_{T}(t,L)=a_{0}\cdot\exp\left(\frac{t}{L}\right)\,.$ (56)
We shall now determine the ratio $t/L$.
How can we interpret this tree? The tree represents the process to change the Wilson loop as encoded into the topology change $S^{3}\rightarrow\Sigma$. This process is driven by the stochastic geometry flow. The possible changes of the Wilson loop, as driven by the noise in the stochastic geometry flow, can be in principle arranged into a tree. There occur to be many small changes, and only a few larger changes. We hence interpret the tree in following sense: the higher levels represent small changes, whilst the lower levels represent the larger changes. From the formal point of view, we introduce intermediate steps between $S^{3}$ and $\Sigma$, denoted by $Y_{n}$. To every $Y_{n}$ we assign the length scale $L$. Then $Y_{\infty}$ must be identified with $\Sigma=Y_{\infty}$. We then remark that $Y_{n}$ are hyperbolic 3-manifolds with a fixed length scale called Mostow-Prasad rigidity — the volume of a hyperbolic 3-manifold is a topological invariant. Then we define
$|^{3}R|=\frac{1}{L^{2}}$
to be the absolute value of the scalar curvature of $Y_{n}$ with respect to the scale $L$. By a simple integration with respect to the metric $h$ of the 3-manifold $Y_{n}$, we obtain
$\frac{1}{L^{2}}=\frac{\intop_{Y_{n}}\,|^{3}R|\sqrt{h}d^{3}x}{\intop_{Y_{n}}\sqrt{h}d^{3}x}$
expressing the constant scalar curvature. In the following, we will relate the expression
$\intop_{\Sigma}tr\left(A\wedge dA+\frac{2}{3}A\wedge A\wedge A\right)=8\pi^{2}CS(\Sigma)\,,$
known as Chern-Simons invariant of a 3-manifold $\Sigma$, to the previous curvature integral. Using ideas of Witten Witten (1988/89, 1989, 1991), we will interpret the connection $A$ as an $ISO(2,1)$ connection. Note that $ISO(2,1)$ is the Lorentz group $SO(3,1)$ by Wigner-Inönü contraction, or the isometry group of the hyperbolic geometry. At this purpose, we choose
$A_{i}=\frac{1}{\ell}e_{i}^{a}P_{a}+\omega_{i}^{a}J_{a}\,,$ (57)
with $\ell$ a length scale and $A=A_{i}dx^{i}$ a 1-form valued in the Lie algebra $ISO(2,1)$. Thus the generators $P_{a}$, $J_{a}$ fulfill the commutation relations
$[J_{a},J_{b}]=\epsilon_{abc}J^{c}\,,\qquad[P_{a},P_{b}]=0\,,\qquad[J_{a},P_{b}]=\epsilon_{abc}P^{c}\,,$
with pairings $\langle J_{a},P_{b}\rangle=Tr(J_{a}P_{a})=\delta_{ab}$, $\langle J_{a},J_{b}\rangle=0=\langle P_{a},P_{b}\rangle$. We then obtain for the curvature
$F_{ij}$ $=$ $\frac{1}{\ell}P_{a}\left(\partial_{i}e_{j}^{a}-\partial_{j}e_{i}^{a}+\epsilon^{abc}(\omega_{ib}e_{jc}+e_{ib}\omega_{jc})\right)+$
$J_{a}\left(\partial_{i}\omega_{j}^{a}-\partial_{j}\omega_{i}^{a}+\epsilon^{abc}\omega_{ib}\omega_{jc}\right)\,.$
In what follows, we move from the expression $A\wedge F$, and use the pairing $\langle\,,\,\rangle=Tr(\,)$ following the MacDowell-Mansouri approach — see MacDowell and Mansouri (1977). Then we obtain
$tr(A\wedge F)=\frac{1}{\ell}e\wedge(d\omega+\omega\wedge\omega)+\frac{1}{\ell}\omega\wedge(de+\omega\wedge e)\,,$
where the second term provides $\omega\wedge T$, with $T$ torsion. The first term encodes the structure $e\wedge R$, with curvature 2-form $R$, which reproduces the scalar curvature ${}^{3}R$ multiplied by the volume form (in the first order formalism) within the Einstein-Hilbert action. Therefore, for vanishing torsion $T=0$, we obtain
$\intop_{\Sigma}tr(A\wedge F)=\frac{1}{\ell}\intop_{\Sigma}\,^{3}R\sqrt{h}d^{3}x\,.$
By performing an auxiliary scaling $\omega\rightarrow\frac{3}{2}\omega$, we obtain the new connection
$\frac{2}{3}A_{i}=\frac{2}{3\ell}e_{i}^{a}P_{a}+\omega_{i}^{a}J_{a}\,,$
and then finally the relation
${\rm Tr}\left(A\wedge(dA+\frac{2}{3}A\wedge A)\right)=\frac{3}{2\ell}e\,\wedge(d\omega+\omega\wedge\omega)\,,$
or
$8\pi^{2}\cdot\ell\cdot CS(\Sigma)=\frac{3}{2}\intop_{\Sigma}\,^{3}R\sqrt{h}d^{3}x\,.$ (58)
In (57) we have been forced to introduce the length scale $\ell$ for the translation generators $P_{a}$. This rescaling can now be parametrised by a coordinate $t$. Therefore, we redefine the expression $8\pi^{2}\ell$ in (58) as
$t\cdot CS(\Sigma_{2})=\frac{3}{2}\intop_{\Sigma_{2}}\,^{3}R\sqrt{h}d^{3}x\,,$ (59)
where the extra factor $8\pi^{2}$ (which equals $4\cdot{\rm vol}(S^{3})$) is the normalisation of the curvature integral. This normalisation of the curvature changes its absolute value into
$|^{3}R_{ren}|=\frac{1}{8\pi^{2}L^{2}}\,.$ (60)
This naturally suggests a choice of the scale factor from its relation with the volume $L=\sqrt[3]{{\rm vol}(Y_{\infty})/(8\pi^{2})}$, where again $Y_{\infty}=\Sigma$. We finally obtain the formal relation
$\intop_{Y_{\infty}}\,|^{3}R_{ren}|\sqrt{h}\,d^{3}x=\intop_{Y_{\infty}}\frac{1}{8\pi^{2}L^{2}}\sqrt{h}\,d^{3}x=L^{3}\cdot\frac{1}{L^{2}}=L\,,$ (61)
by using
$L^{3}=\frac{{\rm vol}(Y_{\infty})}{8\pi^{2}}=\frac{1}{8\pi^{2}}\intop_{Y_{\infty}}\sqrt{h}\,d^{3}x\,,$
which is in agreement with the previous normalisation. We may now use the absolute value of the curvature $|^{3}R|$ and of the Chern-Simons invariant $|CS(\Sigma)|$. By (59) and (61), and using
$\frac{t}{L}=\frac{3}{2\cdot CS(\Sigma)}\,,$
and Eq. (56), we derive the following exponential behavior
$a(t)=a_{0}\cdot e^{t/L}=a_{0}\cdot\exp\left(\frac{3}{2\cdot CS(\Sigma)}\right)\,.$
We may discuss the topology change $S^{3}\rightarrow\Sigma$. The change by adding a single pair of 1-/2-handle produces in the simplest case the so-called Mazur manifold with boundary the Brieskorn sphere $\Sigma=\Sigma(2,5,7)$. Then we will calculate the change in the scale by the above formula. As previously noted, we assume a Planck-size $(L_{P})$ — natural (low-energy) scale for the geometry side — for the 3-sphere $S^{3}$. This changes to the scale $a$ of $\Sigma(2,5,7)$ according to
$a=L_{P}\cdot\exp\left(\frac{3}{2\cdot CS(\Sigma(2,5,7))}\right)\,,$
with the Chern-Simons invariant
$CS(\Sigma(2,5,7))=\frac{9}{4\cdot(2\cdot 5\cdot 7)}=\frac{9}{280}\,.$
This provides a tremendous result: the energy scale, which is order of the Planck scale, i.e. $10^{-34}m$, changes to value $10^{-15}m$. In particualr, we will obtain
$a_{1}=L_{P}\cdot\exp\left(\frac{140}{3}\right)\approx 7.5\cdot 10^{-15}m\,.$
This can be finally related to an energy scale. The estimate of its related Compton length provides the value of the relevant energy scale entering the topology changes processes hat we speculate are at the origin of confinement, namely 165 MeV. This is indeed comparable to the energy scale of QCD, which lies between 217 MeV and 350 MeV — see e.g. Refs. Sekhar Chivukula (2004); Deur et al. (2016).
VIII Proposal for QCD confinement
@@
Geometric fluctuations due to the stochastic geometric flow induce changes in the field lines, following the pattern of knot concordance described in Section VI. Within this framework, the expectation of the knotted and linked Wilson loops is computed by considering the vev of the product of two Wilson loops as
@@
where $S_{\rm YM}$ denotes the Yang–Mills action, and the normalisation constant $Z$ is the partition function. Delving into the analogy with the Chern-Simons case studied in Ref. Cotta-Ramusino et al. (1990), we expect $\langle{W}_{\gamma_{1}}(A){W}_{\gamma_{2}}(A)\rangle$ to depend on the linking of $\gamma_{1}$ and $\gamma_{2}$. Within the Chern-Simons framework, it is known that switching crossings produces certain skein relations — see e.g. as shown in Ref. Cotta-Ramusino et al. (1990). This explains why the vev of Wilson loops provides link invariants, such as the Jones polynomial.
In order to support this conjecture, we show how to estimate the path integral in Eq. (62). At this purpose, we move from Eq. (49), and use the relation in Eq. (48). Then we recast the Wilson functional as
[…]
with the surface $\Gamma_{1}$ such that $\partial\Gamma_{1}=\gamma_{1}$. Schematically, we obtain
[…]
with $F$ denoting the curvature of the connection $A$. The Yang-Mill action consists of the term
$g^{\mu\rho}g^{\nu\lambda}F_{\mu\nu}F_{\rho\lambda}\,,$
which we write schematically as $g^{2}F^{2}$. Thus we obtain the integral
[…]
over the spacetime $M$. To proceed with the calculation, we have to introduce the support function ${\rm supp}(\Gamma_{1})$, with
[…]
The following relations are straightforwardly recovered:
[…]
$({\rm supp}(\Gamma_{1}))^{2}$ $=$ ${\rm supp}(\Gamma_{1})\,.$
We may now recast the Wilson functionals and the Yang-Mills action in one integral over $M$, with kernel
$F^{2}g^{2}+({\rm supp}(\Gamma_{1})+{\rm supp}(\Gamma_{2}))F\,.$
This latter can be further recast according to
$\left(Fg+\frac{({\rm supp}(\Gamma_{1})\!+\!{\rm supp}(\Gamma_{2}))}{2g}\right)^{2}\!-\frac{\left({\rm supp}(\Gamma_{1})+{\rm supp}(\Gamma_{2})\right)^{2}}{(4g^{2})}\,.$
From first expression, which is quadratic in the field strength, we can directly proceed to calculate the path integral. While from the second expression we obtain (up to constants)
[…]
The latter integral can be interpreted as the intersection between the two surfaces $\Gamma_{1}$ and $\Gamma_{2}$. This is equivalent to the linking number $lk(\gamma_{1},\gamma_{2})$ between the knots $\gamma_{1}$ and $\gamma_{2}$, i.e.
[…]
where “$lk(\gamma_{1},\gamma_{2})$” denotes the linking number between the knots $\gamma_{1}$ and $\gamma_{2}$. Consistently with these relations, we derive up to constants the behaviour
$\langle{W}_{\gamma_{1}}(A){W}_{\gamma_{2}}(A)\rangle\sim e^{\imath\,{\rm Area}(\Gamma_{1})+\imath\,{\rm Area}(\Gamma_{2})}\,e^{2\imath\,lk(\gamma_{1},\gamma_{2})}\,,$ (63)
which is our desired result. The details of this calculation will be presented in a forthcoming work.
IX Conclusions and outlook
We outlined the proposal for a novel geometric phase approach to confinement in Yang–Mills theories. The crucial aspects of the framework we have developed can be summarised as follows:
@@ IX Conclusions and outlook
  • ii) the gravitational back-reaction to the Yang–Mills fields induces stochastic fluctuations of topology of the sub-manifolds that represent matter;
  • iii) in the dual picture, changes of topology corresponds to the generation of gauge vortices, which finally source the geometric phase of hadrons’ ground states.
@@
First of all, the formation at the fermi scale of a non-vanishing condensate is eminently related to the topological properties of the manifold. But the energy scale that defines the size of the ground states is still related to dimensional transmutation. Nonetheless, it might be tempting in this framework to relate the energy scale of confinement to the amplitude of the stochastic noises that enter the stochastic Einstein–Yang–Mills flow. In Lulli et al. (2022) the amplitude of the gravitational stochastic noise was traced back to the variance of its fluctuations, and shown to be proportional to the cosmological constant. On the other hand, for the gauge fields it might be natural to relate the origin of the amplitude of the stochastic noise to chaos effects in the dynamics of the chromo-electric and chromo-magnetic fluxes.
@@
Dimensional transmutation can be also reconsidered in the framework we are proposing. We may indeed account for the breakdown of the conformal symmetry at the quantum level, as being induced by the stochastic gauge-geometry flow, while the system is out of equilibrium. The stochastic time regulates the out-of-equilibrium dynamics, making evident that the breakdown of the conformal symmetry (at the quantum level) emerges at non-zero stochastic time distances, and out of equilibrium. Scales can be then individuated in the theory through the coupling between gravity and matter. It is indeed the Newton constant $G$ that provides a dimension-full constant to introduce new physical scale. On the other hand, it is the interplay between the coupling constant of the theory of gravity and the specific coupling constant of matter that generate the physical relevant scale for the QFT under investigation.
Two specific examples of emergence of physical scales through dimensional transmutation — induced by the coupling with matter, and reminiscent of the conformal anomaly as sourced by the trace of the energy-momentum tensor — are provided by the theory of QED (the Bohr radius of the electron, which regulates the “size” of the Hydrogen atom) and the theory of QCD (the $\Lambda_{\rm QCD}$ scale). In these two cases, the physical scales should be provided by a running, induced by the stochastic gauge-geometry flow, of the values of $G$, together with a rescaling with the respect to $g_{\rm em}$ and $g_{\rm s}$ of the gravitational coupling constant $G$. Hence, it would be tempting to consider
@@
having denoted with $\delta^{\rm QCD}\Lambda$ the contribution from the QCD sector to the cosmological constant $\Lambda$.
The so-called “Vacuum Catastrophe” reflects the fundamental fact that the energy density of the quark-gluon condensate responsible for the color confinement in QCD is off by over forty orders of magnitude and has a wrong sign compared to the observable cosmological $\Lambda$-term value. In the Yang–Mills effective action approach Savvidy (1977); Matinyan and Savvidy (1978) it is possible to show that the exact compensation of the electric (metastable) and magnetic (stable) gluon condensate components — localised into vacuum domains spatially separated by domain walls in the QCD vacuum (averaged over space-time volumes above the fermi scale) — is the necessary and sufficient condition for confinement in QCD Addazi et al. (2019a, b). Provided that the energy scales of “electric gluon” and “magnetic gluon” condensation are not the same, the latter are formed at different space-time separations. However, they evolve towards the same values of the energy density, but with opposite signs, due to their cosmological attractor nature. This effectively causes the cancellation of “electric gluon” and “magnetic gluon” contributions to the QCD ground state in macroscopic volumes, i.e. in the deep infrared limit of the theory Pasechnik et al. (2013a); Addazi et al. (2022). The gravitational interactions then break the discrete mirror symmetry of the QCD vacuum already in the semi-classical approximation. Such a breaking causes a non-zeroth (positive) gravitational correction to the resulting effective QCD-induced cosmological constant $\delta^{\rm QCD}\Lambda$, found to be close (within an order of magnitude) to the observed value Pasechnik et al. (2013b).
Such a cancellation emerges in experiment as a complete disappearance of the colored degrees of freedom from the theory in the IR limit, i.e. with characteristic wave lengths beyond the fermi scale, as they are “trapped” within the domains at the QCD length scale. This is fully compatible with the classical limit of the Yang–Mills theory, in which the conformal anomaly dynamically vanishes and only a matter-like (non-relativistic) QCD medium (hadron gas) remains at large separations. For the exact cancellation, and hence for color confinement, the restoration of a discrete (mirror) symmetry between the “electric gluon” and “magnetic gluon” contributions at the level of the effective Lagrangian is an intrinsic property of pure Yang–Mills theory and the RG flow equations.